OpenStudy (anonymous):

The minimum and maximum temperature on a cold day in Lollypop Town can be modeled by the equation below: 2|x − 6| + 14 = 38 What are the minimum and maximum temperatures for this day? x = −9, x = 21 x = −6, x = 18 x = 6, x = −18 No solution

pooja195 (pooja195):

Add -14 to both sides. 2(|x−6|)+14+−14=38+−14 2(|x−6|)=24 Step 2: Divide both sides by 2.

pooja195 (pooja195):

\[\huge~\rm \frac{ 2(|x−6|)=24 }{ 2~~~~=~~~~~~~2 }\]

OpenStudy (anonymous):

I have no idea how to divide like that lol